A rectangle's length is 9 inches greater than its width. If the perimeter of the rectangle is 66 inches, find the length
Answer by waynest(281)

length = w + 9
P = 2l + 2w
66 = 2(w + 9) + 2w
66 = 2w + 18 + 2w
66 - 18 = 4w + 18 - 18
48 = 4w
48/4 = 4w/4
12 = w
check:
2(12) + 2(12 + 9) = 66
24 + 42 = 66
66 = 66